# Kernelized Decoded Quantum Interferometry

### v2

@@ Appendix C Polynomial-phase concentration for OPI
This appendix records a concrete version of the polynomial-phase concentration (PPC) property used in Section IV.1 and connects it to the head-mass quantity $\Sigma_{K}(\ell,\eta;d)$.
@@ C.1 Formal PPC property
Definition C.1 (Polynomial-phase concentration (PPC))
@@ 2. Explicit PPC for quadratic OPI with contiguous evaluation sets
Lemma C.1 (Exact PPC for quadratic OPI over contiguous evaluation sets)
@@
After normalization, the spectrum $\alpha=F_{p}K_{-a}g/\|g\|_{2}$ has $|\alpha_{m^{\star}}|^{2}=1$ and $\alpha_{m}=0$ for $m\neq m^{\star}$, so the entire spectral mass lies in the head of size $d^{\star}=1$ with no tail. This is exactly the PPC property of Definition C.1 with $r=2$, $d^{\star}=1$, and $\varepsilon_{2,p}=0$.
Corollary C.1 (Concrete PPC constants for quadratic OPI)
@@ C.2 Justification via discrete stationary phase
The PPC property can be motivated by discrete stationary-phase and linear canonical transform identities. Quadratic-phase kernels exactly diagonalize quadratic exponential sums and approximately diagonalize higher-degree polynomial phases, with tails controlled by Weyl-type exponential-sum bounds; see, for example, [HealyLCT2016, Rabiner1969CZT] for LCT/fractional Fourier transform identities and [JordanNature2025] for how the OPI shaping map inherits low-degree algebra from the underlying polynomial $h$. A detailed coding of these classical estimates into the DQI context is beyond our scope; Definition C.1 encapsulates the resulting concentration behavior in a form suitable for our head-mass analysis.
Lemma C.2 (PPC for affine OPI instances)
Let $p$ be prime and consider an OPI instance over $\mathbb{F}_{p}$ with evaluation points $\{0,1,\dots,p-1\}$ and a polynomial $f(x)=ax+b$ of degree at most $1$. Suppose the shaping map of Section IV is chosen so that the shaped amplitudes before the kernel are proportional to
@@
Then with the identity kernel $K=\mathbb{I}$ and the pp–ary QFT $F_{p}$, the post–interference spectrum $F_{p}\psi$ is supported on a single mode. In particular, Definition C.1 holds with degree bound $d=1$, head size $k=1$, and tail parameter $\varepsilon_{p}(1,1)=0$.
proof C.2
@@
The inner sum vanishes unless $m\equiv a\pmod{p}$, in which case it equals $p$. Thus all spectral mass is concentrated on the unique index $m^{\star}$ with $m^{\star}\equiv a\pmod{p}$, and the normalized spectrum satisfies $|(F_{p}\psi)_{m^{\star}}|^{2}=1$ and $(F_{p}\psi)_{m}=0$ for $m\neq m^{\star}$. This is exactly the PPC property with $k=1$ and zero tail.
@@ C.3 From PPC to head mass
Lemma C.3 (PPC implies concentrated head mass)
Suppose the kernel family $\{K(\theta)\}$ satisfies PPC (Definition C.1) for degree bound $r$ and head size $d_{\star}$. Then for every admissible OPI instance there exists a parameter $\theta^{\star}$ such that the normalized amplitudes $\alpha=\mathcal{F}K(\theta^{\star})g/\|g\|_{2}$ satisfy
@@
proof C.3
This is a restatement of Definition C.1 with $d=d_{\star}$; no additional argument is required.
@@
where $\underline{\eta}$ is the minimum per-mode attenuation on the head and $\Delta(w)$ is the leakage term defined in Lemma B.1. This is the head-mass estimate used in Theorem IV.1.
@@ Appendix D Block-local alignment for LDPC-like ensembles
We next formalize the block-local alignment (BLA) property used in Section IV.2 and relate it to gains in the head-spectrum mass for sparse Max-XORSAT / LDPC-like instances.
@@ D.2 Block-local alignment property
Definition D.1 (Block-local alignment (BLA))
@@ D.3 From BLA to weighted head mass
Lemma D.1 (BLA implies weighted head-mass gain)
Assume the noise model and standing conditions of Section II.3 and Appendix B, and let $K$ satisfy the BLA property of Definition D.1 with gain $\Delta\Sigma_{\mathrm{loc}}$. Let $\underline{\eta}:=\min_{s\in\mathcal{S}_{d}(\alpha)}\eta_{K}(s)$ be the minimum per-mode attenuation on the head for the kernel $K$. Then for head size $d=\Theta(n/b)$ the noise-weighted head mass satisfies
@@
Step 1: use BLA to compare unweighted head masses. The BLA property (Definition D.1) for the block-local kernel $K$ asserts that there is an additive gain $\Delta\Sigma_{\mathrm{loc}}>0$ such that, for head size $d=\Theta(n/b)$,
@@
where the last inequality uses only the fact that $\Delta(w)\geq 0$. This is exactly the claimed bound.
Combining Lemma D.1 with the noise-aware lower bound in Theorem B.1 and the LDPC decoder response $F_{\mathrm{LDPC}}$ from Section II.3 yields Theorem IV.2 in the main text: block-local kernels that satisfy BLA induce an additive improvement in $\Sigma_{K}$, which in turn shifts BP thresholds and improves the k-DQI approximation guarantees.
@@ 4. Specialization to regular $(3,6)$ LDPC ensembles on the BEC
Proposition D.1 (BLA-induced DE improvement for regular $(3,6)$ ensembles)
@@
Since $0<\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}\leq 1$ by assumption, both quantities are strictly smaller than their threshold counterparts, proving (65)–(66). The fact that the recursion is contractive in a neighbourhood of $x^{\star}$ follows from standard one-dimensional fixed-point theory: a differentiable map with slope strictly less than 1 in modulus is locally attracting. The linear dependence on $\Delta\Sigma_{\mathrm{loc}}$ is explicit in the prefactors.
Remark D.1 (Calibration of $\kappa_{3,6}$ and connection to simulations)
In practice, $\kappa_{3,6}$ is estimated once by matching the head-mass gain $\Delta\Sigma_{\mathrm{loc}}$ to the observed shift of the DE waterfall curve in numerical experiments (see Fig. 2). Proposition D.1 then shows that any positive calibrated value of $\kappa_{3,6}$ suffices to move the DE map strictly below the diagonal at the original threshold point and to dampen the derivative peak. Thus, within the DE surrogate, the BLA-induced gain $\Delta\Sigma_{\mathrm{loc}}$ translates directly into a provable margin of BP stability.
@@ Appendix E Smoothness of the Optimization Landscape
To rigorously justify the efficiency of the variational approach proposed in Section VI, we prove that the objective function $\Sigma_{K}(\bm{\theta})$ is Lipschitz continuous with respect to the kernel parameters. This rules out “shattering” of the landscape, ensuring that gradient-based or gradient-free optimizers can reliable navigate to the optimum.

### v3

@@ Appendix C Polynomial-phase concentration for OPI
This appendix records a concrete version of the polynomial-phase concentration (PPC) property used in Section 2.5 and connects it to the head-mass quantity $\Sigma_{K}(\ell,\eta;d)$.
@@ C.1 Formal PPC property
Definition 1 (Polynomial-phase concentration (PPC)).
@@ 2. Explicit PPC for quadratic OPI with contiguous evaluation sets
Lemma C.1 (Exact PPC for quadratic OPI over contiguous evaluation sets).
@@
After normalization, the spectrum $\alpha=F_{p}K_{-a}g/\|g\|_{2}$ has $|\alpha_{m^{\star}}|^{2}=1$ and $\alpha_{m}=0$ for $m\neq m^{\star}$, so the entire spectral mass lies in the head of size $d^{\star}=1$ with no tail. This is exactly the PPC property of Definition C.1 with $r=2$, $d^{\star}=1$, and $\varepsilon_{2,p}=0$. ∎
Corollary C.1 (Concrete PPC constants for quadratic OPI).
@@ C.2 Justification via discrete stationary phase
The PPC property can be motivated by discrete stationary-phase and linear canonical transform identities. Quadratic-phase kernels exactly diagonalize quadratic exponential sums and approximately diagonalize higher-degree polynomial phases, with tails controlled by Weyl-type exponential-sum bounds; see, for example, [Healy2016LCTBook, Rabiner1969CZT] for LCT/fractional Fourier transform identities and [JordanNature2025] for how the OPI shaping map inherits low-degree algebra from the underlying polynomial $h$. A detailed coding of these classical estimates into the DQI context is beyond our scope; Definition 1 encapsulates the resulting concentration behavior in a form suitable for our head-mass analysis.
Lemma C.2 (PPC for affine OPI instances).
Let $p$ be prime and consider an OPI instance over $\mathbb{F}_{p}$ with evaluation points $\{0,1,\dots,p-1\}$ and a polynomial $f(x)=ax+b$ of degree at most $1$. Suppose the shaping map of Section 2.5 is chosen so that the shaped amplitudes before the kernel are proportional to
@@
Then with the identity kernel $K=\mathbb{I}$ and the pp–ary QFT $F_{p}$, the post–interference spectrum $F_{p}\psi$ is supported on a single mode. In particular, Definition 1 holds with degree bound $d=1$, head size $k=1$, and tail parameter $\varepsilon_{p}(1,1)=0$.
Proof.
@@
The inner sum vanishes unless $m\equiv a\pmod{p}$, in which case it equals $p$. Thus all spectral mass is concentrated on the unique index $m^{\star}$ with $m^{\star}\equiv a\pmod{p}$, and the normalized spectrum satisfies $|(F_{p}\psi)_{m^{\star}}|^{2}=1$ and $(F_{p}\psi)_{m}=0$ for $m\neq m^{\star}$. This is exactly the PPC property with $k=1$ and zero tail. ∎
@@ C.3 From PPC to head mass
Lemma C.3 (PPC implies concentrated head mass).
Suppose the kernel family $\{K(\theta)\}$ satisfies PPC (Definition 1) for degree bound $r$ and head size $d_{\star}$. Then for every admissible OPI instance there exists a parameter $\theta^{\star}$ such that the normalized amplitudes $\alpha=\mathcal{F}K(\theta^{\star})g/\|g\|_{2}$ satisfy
@@
Proof.
This is a restatement of Definition 1 with $d=d_{\star}$; no additional argument is required. ∎
@@
where $\underline{\eta}$ is the minimum per-mode attenuation on the head and $\Delta(w)$ is the leakage term defined in Lemma B.1. This is the head-mass estimate used in Theorem 2.5.
@@ Appendix D Block-local alignment for LDPC-like ensembles
We next formalize the block-local alignment (BLA) property used in Section 2.5 and relate it to gains in the head-spectrum mass for sparse Max-XORSAT / LDPC-like instances.
@@ D.2 Block-local alignment property
Definition 2 (Block-local alignment (BLA)).
@@ D.3 From BLA to weighted head mass
Lemma D.1 (BLA implies weighted head-mass gain).
Assume the noise model and standing conditions of Section 2.3 and Appendix B, and let $K$ satisfy the BLA property of Definition 2 with gain $\Delta\Sigma_{\mathrm{loc}}$. Let $\underline{\eta}:=\min_{s\in\mathcal{S}_{d}(\alpha)}\eta_{K}(s)$ be the minimum per-mode attenuation on the head for the kernel $K$. Then for head size $d=\Theta(n/b)$ the noise-weighted head mass satisfies
@@
Step 1: use BLA to compare unweighted head masses. The BLA property (Definition 2) for the block-local kernel $K$ asserts that there is an additive gain $\Delta\Sigma_{\mathrm{loc}}>0$ such that, for head size $d=\Theta(n/b)$,
@@
where the last inequality uses only the fact that $\Delta(w)\geq 0$. This is exactly the claimed bound. ∎
Combining Lemma D.1 with the noise-aware lower bound in Theorem 4 and the LDPC decoder response $F_{\mathrm{LDPC}}$ from Section 2.3 yields Theorem 2.5 in the main text: block-local kernels that satisfy BLA induce an additive improvement in $\Sigma_{K}$, which in turn shifts BP thresholds and improves the k-DQI approximation guarantees.
@@ 4. Specialization to regular $(3,6)$ LDPC ensembles on the BEC
Proposition 5 (BLA-induced DE improvement for regular $(3,6)$ ensembles).
@@
Since $0<\kappa_{3,6}\,\Delta\Sigma_{\mathrm{loc}}\leq 1$ by assumption, both quantities are strictly smaller than their threshold counterparts, proving (64)–(65). The fact that the recursion is contractive in a neighbourhood of $x^{\star}$ follows from standard one-dimensional fixed-point theory: a differentiable map with slope strictly less than 1 in modulus is locally attracting. The linear dependence on $\Delta\Sigma_{\mathrm{loc}}$ is explicit in the prefactors. ∎
Remark 1 (Calibration of $\kappa_{3,6}$ and connection to simulations).
In practice, $\kappa_{3,6}$ is estimated once by matching the head-mass gain $\Delta\Sigma_{\mathrm{loc}}$ to the observed shift of the DE waterfall curve in numerical experiments (see Fig. 2). Proposition 5 then shows that any positive calibrated value of $\kappa_{3,6}$ suffices to move the DE map strictly below the diagonal at the original threshold point and to dampen the derivative peak. Thus, within the DE surrogate, the BLA-induced gain $\Delta\Sigma_{\mathrm{loc}}$ translates directly into a provable margin of BP stability.
@@ Appendix E Smoothness of the Optimization Landscape
To rigorously justify the efficiency of the variational approach proposed in Section 3.3, we prove that the objective function $\Sigma_{K}(\bm{\theta})$ is Lipschitz continuous with respect to the kernel parameters. This rules out “shattering” of the landscape, ensuring that gradient-based or gradient-free optimizers can reliable navigate to the optimum.